system of simultaneous equations by the addition method.
x+y=1
x-y=-7

whats my solution?
1
Expert's answer
2012-03-13T07:04:09-0400
x+y=1
x-y=-7
Let'sadd the equations:
2x=-6
whence x=-3
from the first equation of the system we get y=1-x
whence y=1-(-3)=4